Statistics students in Oxnard College sampled 11 textbooks in the Condor bookstore and recorded the number of pages in each textbook and its cost. The bivariate data are shown below Number of Pages (Cost(y) 135.72 110.94 114.33 65.89 38.9 55.81 88.42 98.42 124.17 74.46 53.15 944 838 841 453 230 337 634 734 909 442 355 A student calculates a linear model (Please show your answers to two decimal places) Use the model to estimate the cost when number of pages is 534 Cost- places.) (Please show your answer to 2 decimal
Answer #1

a)

Consider the given table

X 944 838 841 453 230 337 634 734 909 442 355 y|135.72 110.94 114.33 65.89 38.9 55.81 88.42 98.42 124.17 74.46 53.15

X Y XY XX
944 135.72 128119.68 891136
838 110.94 92967.72 702244
841 114.33 96151.53 707281
453 65.89 29848.17 205209
230 38.9 8947 52900
337 55.81 18807.97 113569
634 88.42 56058.28 401956
734 98.42 72240.28 538756
909 124.17 112870.53 826281
442 74.46 32911.32 195364
355 53.15 18868.25 126025

Find the sum of column as

\sum{X} = 6717 ~,~ \sum{Y} = 960.21 ~,~ \sum{X \cdot Y} = 667790.73 ~,~ \sum{X^2} = 4760721

\begin{aligned} a &= \frac{\sum{Y} \cdot \sum{X^2} - \sum{X} \cdot \sum{XY} }{n \cdot \sum{X^2} - \left(\sum{X}\right)^2} = \frac{ 960.21 \cdot 4760721 - 6717 \cdot 667790.73}{ 11 \cdot 4760721 - 6717^2} \approx 11.827 \\ \\b &= \frac{ n \cdot \sum{XY} - \sum{X} \cdot \sum{Y}}{n \cdot \sum{X^2} - \left(\sum{X}\right)^2} = \frac{ 11 \cdot 667790.73 - 6717 \cdot 960.21 }{ 11 \cdot 4760721 - \left( 6717 \right)^2} \approx 0.124\end{aligned}

y=bx+a

y=0.12x+11.83

b)

Substitute x= 534 to obtain

y= 0.12 *534 +11.83

y=$75.91

Please let me know in comment if you have any doubt. Thanks